OpenStudy (anonymous):

use the product to sum formulas to write the product as a sum or difference 10 cos 75° cos 15°

OpenStudy (campbell_st):

its basically 5cos(75 + 15) + 5 cos(75 - 15) the terms 5sin75sin15 -5sin75sin15 = 0 will appear

OpenStudy (anonymous):

thanks!

OpenStudy (anonymous):

but where did you get the 5 from?

OpenStudy (campbell_st):

its the coefficient of cos(75 + 15) so 5cos(75 + 15) = 5(cos(75)cos(15) - sin(75)sin(15)) = 5cos(75)(cos(15) - 5sin(75)sin(15) hope that helps

OpenStudy (anonymous):

I got it. haha It came from the two that was divided into the problem